the density of a liquid that has mass of 7.5E3 mg and a volume of 22.01mL??
Answer #1

Density=Mass/Vol

In this problem Mass=7.5E3mg=7.5 gm

Volume=22.01 ml

Density of Liquid=7.5/22.01=0.34075 gm/ml